# Mitigating Graph Covariate Shift via Score-based Out-of-distribution Augmentation

### v1

@@ 3 Problem Formulation
where $\ell(\cdot,\cdot)$ denotes the loss function for graph classification and the expectation is with respect to graphs under all possible environments. However in practice, the training environments $\mathcal{E}_{\mathrm{tr}}$ may not cover all environments, causing degraded classification performance when applying the learned classifier in unseen test environments. This covariate shift calls for an effective manner to sufficiently explore unseen data distribution or environments during model training. We summarize and discuss the various types of graph covariate shifts in detail in Appendix A.2.
Issues in Graph Augmentation via Environmental Exploration. To augment the training distribution for mitigating graph covariate shift, existing solutions often approach Eq. (1) by separating and augmenting the environments: ${\min_{f}}\ \mathbb{E}_{e\in\{\mathcal{E}_{\mathrm{tr}}\cup\mathcal{E}_{\mathrm{aug}}\}}\mathbb{E}_{(G^{e},Y^{e})\sim P_{e}(G,Y)}[\ell(f(G^{e}),Y^{e})],$ where the augmented environments $\mathcal{E}_{\mathrm{aug}}$ are obtained based on either interpolating explicitly given environmental labels (Li et al., 2024) or perturbing implicitly separated environmental components (Sui et al., 2024; Wu et al., 2022; Miao et al., 2022; Chen et al., 2022). Obtaining accurate environmental labels and components itself could be high-cost and nontrivial tasks, and separating environmental components could be inherently unfeasible (Chen et al., 2024), which limit the practicability of such strategy. In addition, the subgraph perturbations based augmentation is mainly operated by edge dropping (Sui et al., 2024; Rong et al., 2019) and mixup (Han et al., 2022), which is confined to existing subgraphs in training data and could cause invalid samples (e.g., generating molecules that are chemically invalid).
Distribution Augmentation with OOD Control. This work overcomes these limitations by formulating the augmentation problem as a generation-based graph OOD augmentation strategy, which directly models and augments the training distribution, without explicitly requiring the knowledge or separation of environmental information. Specifically, we target on synthesizing an augmented training distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$, which is combined with the original training distribution to obtain the classifier, stated as:
@@
The augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ can be implemented in multiple ways, but it needs to satisfy two principles: (1) $\tilde{P}_{\mathrm{tr}}(G,Y)$ should deviate from $P_{\mathrm{tr}}(G,Y)$ in a controlled manner, and (2) the explored graphs in $\tilde{P}_{\mathrm{tr}}(G,Y)$ should preserve the stable patterns of graphs in $P_{\mathrm{tr}}(G,Y)$. However, current graph generation models (Jo et al., 2022; Martinkus et al., 2022; Vignac et al., 2022) cannot directly generate graphs that meet these two criteria. To address this, we propose a novel score-based generative model in Section 4 that captures the augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ while adhering to both principles.
4 Score-based Out-of-distribution Graph Augmentation
In this section, we present the novel score-based graph augmentation framework, OODA, designed to generate augmented graphs that retain predictable stable features while also exploring new environments. We begin by discussing the score-based generative model for unlabeled graphs and then extend the model to handle out-of-distribution scenarios with controlled adaptation. Thereafter, we illustrate the working principles and implementation details of OODA.
Motivation From the perspective of graph generation, the goal of exploring the training distribution $P_{\mathrm{tr}}(G,Y)$ is to generate OOD graph samples from the conditional distribution $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}\right)$ where $\mathbf{y}_{\text{ood }}$ represents the OOD exploration condition. We assume an exploration variable $\lambda$ controls the extent of exploration within the training distribution $P_{\mathrm{tr}}(G,Y)$. The augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ is then modelled by the conditional graph distribution $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)$, which can be decomposed as follows:
$P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)\propto p\left(G\right)p\left(Y\mid G\right)p\left(\mathbf{y}_{\text{ood }}=\lambda\mid G,Y\right)$ (3)
Existing graph generation models (Jo et al., 2022; Martinkus et al., 2022; Vignac et al., 2022) cannot directly sample graphs from the conditional distribution in Equation (3), as it is infeasible to enumerate all possible $\lambda$ values and their corresponding graphs and labels to compute the normalized probabilities. To overcome this limitation, we propose a novel guidance scheme to direct the score-based generative model (Song et al., 2020) towards the target distribution.
@@ 4 Score-based Out-of-distribution Graph Augmentation
Score-based Graph Generation $p\left(G\right)$ in Equation 3 is the distribution of unlabelled graphs, which can be captured by score-based generative model (Song et al., 2020). The foundational work by (Song et al., 2020) introduced a method for modeling the diffusion process of data into noise and vice versa using stochastic differential equations (SDEs). For graph generation, this diffusion process gradually corrupts graphs into a prior distribution like the normal distribution. The model subsequently samples noise from the prior distribution and learns a score function to denoise the perturbed graphs. Given an unlabeled graph $G$, we use continuous time $t\in[0,T]$ to index the diffusion steps $\left\{G_{t}\right\}_{t=1}^{T}$ of the graph, where $G_{0}$ represents the original distribution and $G_{T}$ follows a prior distribution. The forward diffusion process from the graph to the prior distribution is defined through an Itô SDE:
$\mathrm{d}G_{t}=\mathbf{f}_{t}\left(G_{t}\right)\mathrm{d}t+g_{t}\mathrm{~{}d}\mathbf{w},$ (4)
@@
that incorporates linear drift coefficient $\mathbf{f}_{t}(\cdot):\mathcal{G}\rightarrow\mathcal{G}^{1}$ and scalar diffusion coefficient $g_{t}:\mathcal{G}\rightarrow\mathbb{R}$ related to the amount of noise corrupting the unlabelled graph at each infinitesimal step $t$, along with a standard Wiener process $\mathbf{w}$. In contrast, the reverse diffusion employs SDE that factors in the gradient fields or scores of the perturbed graphs $G_{t}$ for denoising and graph generation from $T$ to 00:
$\mathrm{d}G_{t}=\left[\mathbf{f}_{t}\left(G_{t}\right)-g_{t}^{2}\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)\right]\mathrm{d}\bar{t}+g_{t}\mathrm{~{}d}\overline{\mathbf{w}},$
where $p_{t}\left(G_{t}\right)$ denotes the marginal distribution at time $t$ in forward diffusion, with $\mathbf{f}_{t}\left(G_{t}\right)$ and $g_{t}$ representing the drift and diffusion coefficients, respectively. $\mathrm{~{}d}\overline{\mathbf{w}}$ here is the reverse-time standard Wiener process, and $\mathrm{d}\bar{t}$ is an infinitesimal negative time step. The score network $\bm{s}_{\theta,t}\left(G_{t}\right)$ is trained to approximate the unknown score function $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$. Although the score-based generative model can capture the distribution of unlabeled graphs, it cannot generate the pairs $(G,Y)$ from the OOD distribution. To address this limitation, we introduce a novel OOD guidance scheme designed to generate OOD graphs and their corresponding label from a score-based generative model trained on unlabeled graphs.
@@
Figure 1: Overview of OODA. The diffusion process iteratively transforms an unlabeled original graph $G_{0}$ into noise $G_{T}$. During the denoising process, $G_{t-1}$ is computed using the conditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$, constrained by the target class $\mathbf{y}_{G}$ and the exploration parameter $\lambda$. Ultimately, the clean OOD graph $\tilde{G}_{0}$ is generated.
Score-based Graph Generation with OOD Control To explore the training distribution in a controlled manner, we propose a novel OOD-controlled score-based graph generative model capable of generating OOD graph samples and their corresponding labels. The extent of exploration in the generative process is regulated by the hyperparameter $\lambda$. An overview is provided in Figure 1. Our approach involves sampling $(G,Y)$ from $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)$ and solving the conditional reverse-time SDE:
$\mathrm{d}G_{t}=\left[\mathbf{f}_{t}\left(G_{t}\right)-g_{t}^{2}\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)\right]\mathrm{d}\bar{t}+g_{t}\mathrm{~{}d}\overline{\mathbf{w}}$ (5)
Where $\mathbf{y}_{G}$ denotes the graph’s label and $\mathbf{y}_{\text{ood }}$ specifies the amount of OOD exploration. To sample explored graph instances from $P_{\mathrm{tr}}\left(G,Y\mid\mathbf{y}_{\text{ood }}=\lambda\right)$ using a diffusion model, we note that $\nabla_{G_{t}}\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathbf{y}_{\text{ood }}=\lambda\right)=\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$. The proof is in Appendix A.1. Therefore, the desired conditional reverse-time SDE becomes:
$\mathrm{d}G_{t}=\left[\mathbf{f}_{t}\left(G_{t}\right)-g_{t}^{2}\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)\right]\mathrm{d}\bar{t}+g_{t}\mathrm{~{}d}\overline{\mathbf{w}}$ (6)
According to Equation 3, the conditional score function $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)$ is the sum of three components:
$\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathbf{y}_{\text{ood }}=\lambda\right)=$ $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$ (7)
$+\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$
$+\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$
Based on Bazhenov et al. (2022); Wu et al. (2023), OOD graphs are those with low likelihood under the original distribution $P_{\mathrm{tr}}(G,Y)$. For instance, in the Motif training dataset (Gui et al., 2022), a house motif (stable subgraph) is only connected with wheel graphs, tree graphs or ladder graphs (environmental subgraphs). Consequently, when environmental graphs are explored, the entire graph patterns exist in the low-density region of the original distribution. Inspired by Lee et al. (2023), we model the distribution $p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$ as proportional to the negative exponent of the joint density of $G_{t}$ and $\mathbf{y}_{G}$, $p_{t}\left(G_{t},\mathbf{y}_{G}\right)$:
@@
$p_{t}\left(\mathbf{y}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)\propto p_{t}\left(G_{t},\mathbf{y}_{G}\right)^{-\sqrt{\lambda}}=p_{t}\left(G_{t}\right)^{-\sqrt{\lambda}}p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)^{-\sqrt{\lambda}}$ (8)
Accordingly, the gradient of the log probability for conditional reverse diffusion is expressed as:

### v2

@@ 4. Score-based Conditional Out-of-Distribution Diffusion Augmentation (CODA)
(4) $P_{\mathrm{tr}}\left(G,Y\mid\mathcal{E}_{\text{ood }}=\lambda\right)\propto p\left(G\right)p\left(Y\mid G\right)p\left(\mathcal{E}_{\text{ood }}=\lambda\mid G,Y\right).$
@@
4.1. Preliminaries
Causal View. Following (wu2022discovering), we assume that each input graph $G$ can be decomposed into a causally relevant component $C$ and a spurious (environment-specific) component $S$, e.g., the House motif vs. the Tree base in GOOD-Motif (gui2022good). The label $Y$ depends only on $C$, while $S$ captures variations across environments. These qualitative causal assumptions (pearl2016causal; neuberg2003causality) can be summarized as:
  • $C\to G\leftarrow S$  ($G$ consists of $C$ and $S$).
  • $C\to Y$  ($C$ is sufficient to determine $Y$).
  • $C\not\!\perp\!\!\!\perp S$  ($C$ and $S$ may be statistically dependent, e.g., via latent confounding).
  • $S\perp\!\!\!\perp Y\mid C$  (given $C$, $S$ carries no more information about $Y$).
Low-Density Regions. Class-conditional densities are typically long-tailed (sehwag2022generating), with low-density neighborhoods on the data manifold containing few or no training samples. OOD graphs $G_{\text{ood}}$ tend to reside in these regions (lee2023exploring), and diffusion models are observed to interpolate rather than memorize in such regions (sehwag2022generating). In CODA, we introduce an exploration parameter $\lambda$ to control the extent to which generation is encouraged toward these sparse regions.
Let $\tilde{P}_{e}$ denote the generator’s joint distribution at exploration level $e\equiv\mathcal{E}_{\text{ood}}(\lambda)$. When exploration is disabled ($e=0$), the generator should reproduce the in-distribution data:
$\tilde{P}_{0}(G,Y)\approx P_{\text{in}}(G,Y).$
As the exploration level increases, the generator should allocate more probability mass to OOD neighborhoods of the data manifold. Writing $\mathcal{G}_{\text{ood}}$ for a measurable subset of low-density OOD regions and taking $e_{2}>e_{1}\geq 0$, we expect
$\tilde{P}_{e_{2}}\!\big(G\in\mathcal{G}_{\text{ood}},Y\big)\;>\;\tilde{P}_{e_{1}}\!\big(G\in\mathcal{G}_{\text{ood}},Y\big),$
that is, in shorthand,
$\tilde{P}(G_{\text{ood}},Y\mid\mathcal{E}_{\text{ood}}>0)\;>\;\tilde{P}(G_{\text{ood}},Y\mid\mathcal{E}_{\text{ood}}=0).$
4.2. Forward Diffusion Process
@@
In this work, we focus on diffusion over continuous features to address distribution shifts that occur not only in discrete graph structures but also in continuous node and edge features. The foundational work (song2020score) introduced a method for modeling the diffusion process of data into noise and vice versa using stochastic differential equations (SDEs). For graph generation, this diffusion process gradually corrupts graphs into a prior distribution like the normal distribution. As illustrated in Figure 1, given an unlabeled graph $G$, we use continuous time $t\in[0,T]$ to index the diffusion steps $\left\{G_{t}\right\}_{t=1}^{T}$ of the graph, where $G_{0}$ represents the original distribution and $G_{T}$ follows a prior distribution. The forward diffusion process from the graph to the prior distribution is defined via an Itô SDE:
(5) $\mathrm{d}G_{t}=\mathbf{f}_{t}\left(G_{t}\right)\mathrm{d}t+g_{t}\mathrm{~d}\mathbf{w},$
@@
which incorporates linear drift coefficient $\mathbf{f}_{t}(\cdot):\mathcal{G}\rightarrow\mathcal{G}$ and scalar diffusion coefficient $g_{t}:\mathcal{G}\rightarrow\mathbb{R}$ related to the amount of noise corrupting the unlabelled graph at each infinitesimal step $t$, along with a standard Wiener process $\mathbf{w}$.
@@
Figure 1. The diffusion and reverse processes of OODA. The diffusion process iteratively transforms an unlabeled original graph $G_{0}$ into noise $G_{T}$. During the denoising process, $\tilde{G}_{t-1}$ is computed using the conditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$, constrained by the target class $\mathbf{y}_{G}$ and the exploration parameter $\lambda$. Ultimately, the clean OOD graph $\tilde{G}_{0}$ is generated.
4.3. Score-based Conditional Graph Generation with OOD Control
To steer generation toward controlled OOD characteristics, we condition the reverse-time SDE on the target class $\mathbf{y}_{G}$ and an exploration knob $\lambda\in[0,1]$ that modulates OOD-ness:
(6) $\mathrm{d}G_{t}=\Big[\mathbf{f}_{t}(G_{t})-g_{t}^{2}\,\nabla_{G_{t}}\log p_{t}\!\big(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\mathrm{ood}}=\lambda\big)\Big]\mathrm{d}\bar{t}\;+\;g_{t}\,\mathrm{d}\overline{\mathbf{w}}.$
Since $\mathbf{y}_{G}$ and $\mathcal{E}_{\mathrm{ood}}$ are constants w.r.t. $G_{t}$ (detailed proof in Appendix A.1),
(7) $\nabla_{G_{t}}\log p_{t}\!\big(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\mathrm{ood}}=\lambda\big)=\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big).$
Thus the conditional reverse-time SDE becomes
(8) $\mathrm{d}G_{t}=\Big[\mathbf{f}_{t}(G_{t})-g_{t}^{2}\,\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big)\Big]\mathrm{d}\bar{t}\;+\;g_{t}\,\mathrm{d}\overline{\mathbf{w}}.$
Score Decomposition. By Bayes’ rule,
(9) $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
$=\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)+\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$
$+\nabla_{G_{t}}\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right).$
where the last term, the OOD-ness score, is intractable to model directly. This is because modeling the conditional distribution,
$p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$
would require an infeasible enumeration over a continuous OOD space and all corresponding graphs and labels.
A Tractable OOD Surrogate. From the causal view in §4.1, OOD graphs reside in low-density regions of the training joint $P_{\mathrm{tr}}(G,Y)$. We encode this by assigning higher OOD probability to lower joint density and posit the surrogate
@@ 4.3. Score-based Conditional Graph Generation with OOD Control
(10) $p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$ $\propto p_{t}\left(G_{t},\mathbf{y}_{G}\right)^{-\sqrt{\lambda}}$
$=\big[p_{t}(G_{t})\,p_{t}(\mathbf{y}_{G}\mid G_{t})\big]^{-\sqrt{\lambda}}.$
which yields
(11) $\nabla_{G_{t}}\log p_{t}\!\big(\mathcal{E}_{\mathrm{ood}}=\lambda\mid G_{t},\mathbf{y}_{G}\big)=$
$-\sqrt{\lambda}\,\Big(\nabla_{G_{t}}\log p_{t}(G_{t})+\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})\Big).$
Plugging into (9) gives a simple and tractable conditional score:
(12) $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)=$
$(1-\sqrt{\lambda})\Big(\nabla_{G_{t}}\log p_{t}(G_{t})+\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})\Big).$
We use $0\leq\lambda\leq 1$ so that the guidance stays nonnegative and smoothly trades off between in-distribution sampling ($\lambda=0$) and exploration toward low-density regions ($\lambda\to 1$).
Learning Components. The tractable conditional score in Equation (12) necessitates estimating two key components: the unconditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$ and the class-conditional probability $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$. To this end, we first train a score network $\boldsymbol{s}_{\theta,t}\left(G_{t}\right)$ which takes the noisy graph $G_{t}$ and time $t$ as inputs to approximate the unconditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$. Concurrently, we train a time-dependent classifier $\phi_{t}(G_{t})$ to approximate the conditional class distribution $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$ using inputs $G_{t}$ and $t$. The classifier is trained on noisy pairs $(G_{t},\mathbf{y}_{G})$ to capture time-dependent feature dynamics, rather than on clean $(G_{0},\mathbf{y}_{G})$.
4.4. Working Principles of CODA
Substituting (12) into (8) defines the full sampler. The single hyperparameter $\lambda$ controls the amount of OOD exploration while retaining the class semantics via $\mathbf{y}_{G}$.
Assumptions (from §4.1). Each graph $G$ decomposes into $C$ (causal) and $S$ (non-causal/spurious), with $C\to Y$, $C\not\!\perp\!\!\!\perp S$, and $S\perp\!\!\!\perp Y\mid C$. In particular, $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$.
Proposition 4.1 (Causal structure preservation).
For $0\leq\lambda\leq 1$, the generation induced by (8)–(12) preserves the causal relation $C\!\to\!Y$ under the generated joint $\tilde{P}_{\lambda}(G,Y)$; equivalently, $\tilde{P}_{\lambda}(Y\mid G)=\tilde{P}_{\lambda}(Y\mid C)$ for all $\lambda\in[0,1]$.
Proof. Assumptions (cf. Section 4.1). Each graph decomposes as $G=(C,S)$ with $S\perp\!\!\!\perp Y\mid C$ and hence $P(Y\mid G)=P(Y\mid C)$. The time-dependent classifier is trained on noisy pairs so that $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$, implying $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$.
(i) Vector-Field Structure. From (12),
(13) $\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big)=$